Fix error; remove statement that is not mathematically correct
Line 1:
Line 1:
−
The '''probability''' of an event is a number between 0 (representing impossibility) and 1 (representing certainty), and may be expressed as a simple fraction, decimal fraction, or percentage. When there are a finite number of equally likely outcomes, the probability of an event is given by
+
The '''probability''' of an event is a number between 0 and 1 representing the likelihood of the event's occurrence. It may be expressed as a simple fraction, decimal fraction, or percentage. When there are a finite number of equally likely outcomes, the probability of an event is given by
<math>
<math>
Line 6:
Line 6:
</math>.
</math>.
−
For example, the probability of a tossed fair coin coming up heads is ½ = 0.5 = 50% because there is one favorable outcome (heads) out of 2 possible outcomes (heads and tails). Another common example is standard deck of 52 playing cards: there are 13 cards in each suit: spades, hearts, clubs and diamonds. After the deck is shuffled thoroughly, the probability that the top card on the deck is a heart is <math>13 / 52 = 1 / 4 = 0.20</math>, or 20%.
+
For example, the probability of a tossed fair coin coming up heads is ½ = 0.5 = 50% because there is one favorable outcome (heads) out of 2 possible outcomes (heads and tails). Another common example is standard deck of 52 playing cards: there are 13 cards in each suit: spades, hearts, clubs and diamonds. After the deck is shuffled thoroughly, the probability that the top card on the deck is a heart is <math>13 / 52 = 1 / 4 = 0.25</math>, or 25%.